A building has a ramp to its front doors to accommodate the handicapped. If the distance from the building to the end of the ramp is 25 feet and the height from the ground to the front doors is 7 feet, how long is the ramp? (Round to the nearest tenth.) Answers are below please state which one it is?

24.0 ft

9.9 ft

26.0 ft

5.7 ft

Answer by Boreal(15235)

Draw this right triangle. The answer is the hypotenuse
Pythagorean identity 7^2+25^2=x^2=674
sqrt (674)=25.96 and to nearest tenth 26.0 ft.